The product of two consecutive integers is 2less than the square of the lesser integer. Find the greater of the two integers.
The greater of the two integers is?

Answer by ankor@dixie-net.com(22740)

The two consecutive integers: x, (x+1)
:
The product of two consecutive integers is 2less than the square of the lesser integer.
x(x+1) = x^2 - 2
x^2 + x = x^2 - 2
x^2 - x^2 + x = -2
x = -2
Find the greater of the two integers.
-2 + 1 = -1, is the greater